OpenStudy (anonymous):

A swimming pool has its boundary coinciding with curve of the equation x^(2)+y^(2)-2x+4y-20=0 in meters. How much volume of water is needed to fill the pool top a depth of 1.5 m?

OpenStudy (er.mohd.amir):

(x*x-2*x+1)+(y*y+4*y+4)-1-4-20=0 (x-1)^2+(y+2)^2=25 this represent a circle with cente (1,-2) and radius=5

OpenStudy (er.mohd.amir):

now you know pool is form of sphere use formule of sphere for VOLUME

OpenStudy (phi):

I would guess the pool has the shape of a cylinder, with r=5 and h= 1.5

OpenStudy (anonymous):

oh sorry I fell asleep thak you :)